Question 4 Answer all parts of this question. The postal service of Caicos, an island in the Caribbean, obtains a significantCan someone solve these points for me?
Answer #1

Answer-1:

Unit sales Sales Cost of goods sold Contribution margin f 16.00 £ 17.00 Price Price 110,000 88,000 | £ 1,760,000 £1,496,000 1

The postal service of Caicos make more profit selling souvenir sheets for £ 16.00

Answer-2:

Price elasticity of demand(Ed) = [ln (1 + % change in quantity sold)] / [ln (1 + % change in price)]

= [ln (1 + (- 0.20))] / [ln (1 + (0.0625))]

= [ln (0.80)] / [ln (1.0625)]

= -0.09691 / 0.02633

= -3.68

Answer-3:

Profit-maximizing price = (1 + Profit-maximizing markup on variable cost) × Variable cost per unit

Profit-maximizing markup on variable cost = -1 / (1 + Ed)

= -1 / (1 + (-3.68))

= -1 / -2.68 = 0.3731

Profit-maximizing price = (1 + 0.3731) × £ 1.50 = 1.3731 × £ 1.50 = £ 2.06